Home‎ > ‎Reports‎ > ‎Crew102‎ > ‎

day02


Crew 102 Log Book for March 27, 2011

Photos of the day





Commander's Check-In Report


Commander's Report

Franco Carbognani Reporting

Some of the scientific activities have started. See the Outreach Projects report below.  Whole morning today was spent in making an extra cleanup of the toilet, shower room and closet room. We tried to get in touch with Don Lusco about the propane level getting low.

The interaction with Joshua has been initiated in order to get the grey water system functional. The grey water is needed for one of the experiments planned in order to grow algae. The crew has been in good health and mood (especially after a good plate of spaghetti we had for dinner)


Engineering Report

Franco Carbognani Reporting

Generator/Electricity:

Kitty (Diesel Generator) (Used? Yes/No): Yes

Honey (Gas Generator) (Used? Yes/No): No

Battery Charge Level: (Please let me know where the battery charge indicator is, this info was not passed to me)

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

The laminar sheet of plastic put in place yesterday in order to protect Kitty from the rain has been removed (risk to be blowed away by wind). With the Engineering area roof not in place, all the equipment may suffer during heavy rain.


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0.75

Propane Tank (Full/0.75/0.5/0.25/Empty): 15%

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 0.75

Notes/Comments:

We tried to contact Don to fill up the propane tank but could not get in touch with him.


Oil:

Kitty Oil Quantity (# of Quarts in storage): Same as Day 1

Honey Oil Quantity (# of Quarts in storage): Same as Day 1

ATV Oil Quantity (# of Quarts in storage): Same as Day 1

Car Oil Quantity (# of Quarts in storage): Same as Day 1


Water Status:

Outside Potable Water Tank Level (inches from bottom): 25

Trailer Potable Water Tank Level (inches from bottom): Full

Hab Potable Water Tank Level (inches from bottom): 25

GreenHab Meter Reading: Not Functioning

Notes/Comments:


GreenHab:

Grey water system is not functional at the moment. Verified that the availability 2 bags of 50 pounds Play Sand. Verified that both thermostats are set to 75 degree.


Observatory:

Telescope Used? (Yes/No): No

Observatory Notes/Comments:

Telescope material has been received at the Hollow Mountain. The Crew Astronomer, Yuval Brodsky is in touch with Peter Detterline (The Observatory Focal) in order to set the telescope into operation.


Transportation:

Hab Car (NOT RENTAL)

Used (yes/no): yes

Oil Checked (yes/no):no

Tire Status: good

Odometer at end of day: 123770

Notes/Comments on Hab Car:


ATV 1 (Opportunity):

Used: n

Oil Checked: n

Fuel Consumed:

Tire Status:

Comments:


ATV 2 (Spirit):

Used: n

Oil Checked: n

Fuel Consumed:

Tire Status:

Comments:


ATV 3 (Viking I):

Used: n

Oil Checked: n

Fuel Consumed:

Tire Status:

Comments:


ATV 4 (Viking II):

Used: y/n

Oil Checked: n

Fuel Consumed:

Tire Status:

Comments:


Heating and Ventilation:

Thermostat setting upstairs (Farenheit): 70

Actual temperature upstairs (Farenheit): 69

Thermostat setting downstairs (Farenheit):55

Actual temperature downstairs (Farenheit): 61


Computers/Networking Infrastructure:

Notes/comments on internet/computers?:


General Engineering Notes/Comments:


Health and Safety Officer's Report

Dr. John E. Deaton Reporting

As of this date, the following tasks have been completed:

•       Turnover from prior crew’s (#101) Health and Safety Officer on March 26

•       Inventory and inspection of all first aid supplies on March 26

•       Preparing safety brief to crew on March 26

•       Successful completion of test message to flight surgeon on March 26 upon arriving at MDRS

•       Walk around MDRS to identify any potential safety issues

•       All personnel report satisfactory health status as of this date and time

•       Safety brief provided to all crew on March 27, 2007


Science Report (Human Factors/Psychology)

Dr. John E. Deaton Reporting

Psychology tests (Stress, Leader Evaluation, Exploration Growth Scale, and Emotional Scale) were initiated and will be repeated mid-mission and at the end.  Human performance tests were likewise administered and consisted of multiple tests to measure performance and cognitive processing.

These tests will be repeated every other day until the completion of the mission.


Science Report (Biology)

Lara Vimericati Reporting

The first science project was started today. The main objective was to test resistance to extremely cold temperatures of tardigrades, a phylum of organisms that is known to cope with a complex interplay of stressful conditions.  Their hardy nature makes them very useful to The first science project was started today. The main objective was to test resistance to extremely cold temperatures of tardigrades, a phylum of organisms that is known to cope with a complex interplay of test the resistance of life in space.

After checking for viability of the organisms under the microscope, samples on the slides were placed on dry ice for different amounts of time and their viability was checked upon retrieval. Viability was checked every few minutes up to 1 hour.

Preliminary results showed that tardigrades in hydrated conditions did not display resistance to the very cold temperatures they were subjected when in contact with dry ice. Another test will be run to test whether their resistance to cold temperature increases when they enter an anhydrobiotic state.


Science Report (Geology)

Kavya Manyapu Reporting

The outreach project on geology was conducted today.  The goal of the project was to see how far below the surface of the soil does dry ice have to be without changing from a solid to a gaseous state. The

sublimation of Dry Ice at various depths was tested.

An EVA was conducted for this project. The surface area of the sample of dry ice was determined. The dry ice was place in a shallow depression (about 6 inches) in the soil. The hole was filled with 1 inch of sand at every trial. Each trial was monitored for tree 30 second intervals for sublimation until the fumes were no longer seen. At around depths of 10inches, the dry ice fumes were not visible anymore.

However, there could not be very definitive conclusion drawn within one day of this experiment.  The soil saturation is believed to affect the escape of the dry ice vapors. Since the area around the EVA trail had rain showers this morning, the soil might be well saturated. Hence a repeat of this experiment will be done during the following days of the mission.


Journalist Report

Kavya Manyapu Reporting

Day 2 started out with excitement as the crew could get some rest the previous night. All crew were in excellent health. The day started out with planning the priorities and the experiments for the day.  The commander took care of the non working toilet and cleaned up the hab.

The Human Factors Engineer, John Deaton started his psychological and human performance experiments.  Two outreach projects on behalf of Space Florida were conducted today. The projects involved biology and geology experiments.  The crew biologist, Lara Vimericati conducted the biology experiment of the day. The survival of the phylum of tardigrades in extreme cold environments was tested.

This experiment was designed to explore the possibility of introducing live organisms in the silica soil of Mars.  An EVA was performed by Crew Engineer, Kavya Manyapu to conduct the geology experiment. The ability of silica to insulate the sublimation of Dry Ice was experimented. The subsurface depth below which dry ice does not sublimate was required to be tested. It is believed that this will enable to further the exploration of life by investigating the properties of frozen CO2 and the impact on Martian soil.  The crew Astronomer, Yuval Brodsky, tested the optical astronomy equipment, prepared the observatory for taking observations starting tomorrow.